# Optimal low-rank approximations of posteriors for linear Gaussian inverse problems on Hilbert spaces

### v1

@@ Abstract
For linear inverse problems with Gaussian priors and observation noise, the posterior is Gaussian, with mean and covariance determined by the conditioning formula. We analyse measure approximation problems of finding the best approximation to the posterior in a family of Gaussians with approximate covariance or approximate mean, for Hilbert parameter spaces and finite-dimensional observations. We quantify the error of the approximating Gaussian either with the Kullback–Leibler divergence or the family of Rényi divergences. Using the Feldman–Hajek theorem and recent results on reduced-rank operator approximations, we identify optimal solutions to these measure approximation problems. Our results extend those of Spantini et al. (SIAM J. Sci. Comput. 2015) to Hilbertian parameter spaces. In addition, our results show that the posterior differs from the prior only on a subspace of dimension equal to the rank of the Hessian of the negative log-likelihood, and that this subspace is a subspace of the Cameron–Martin space of the prior.
Keywords: nonparametric linear Bayesian inverse problems, Gaussian measures, low-rank operator approximation, generalised operator eigendecomposition MSC codes: 28C20, 47A58, 60G15, 62F15, 62G05
@@ 1 Introduction
$m_{\textup{pos}}=m_{\textup{pr}}+\mathcal{C}_{\textup{pr}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}(y-Gm_{\textup{pr}}),\quad\mathcal{C}_{\textup{pos}}=\mathcal{C}_{\textup{pr}}-\mathcal{C}_{\textup{pr}}G^{*}(\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{pr}}G^{*})^{-1}G\mathcal{C}_{\textup{pr}},$
see e.g. [31, Example 6.23].
@@
Computational studies of Bayesian inverse problems on high- but finite-dimensional parameter spaces show that the data is often ‘informative’ — i.e., that the posterior differs from the prior — only on a subspace of much lower dimension; see e.g. [14]. Since the posterior is obtained by reweighting the prior by the likelihood, this data-informed subspace is determined by how the concentration of the likelihood interacts with the concentration of prior. In the case of linear Gaussian inverse problems, the concentration of the likelihood and the concentration of the prior are described by the eigenpairs of the Hessian of the negative log-likelihood and the eigenpairs of the prior precision. These ideas were used in [14] to identify low-rank approximations of the posterior covariance matrix. In [30], the optimality of these posterior covariance approximations with respect to a family of spectral loss functions was shown. In particular, the leading generalised eigenvectors of the Hessian-prior precision matrix pencil build a hierarchy of nested low-dimensional subspaces on which the posterior differs from the prior.
@@
The primary goal of this paper is to develop a mathematical analysis of the Bayesian prior-to-posterior update for linear inverse problems where the parameter space is a separable Hilbert space and the observation space is finite-dimensional, using low-rank measure approximation problems to identify optimal Gaussian approximations to the true posterior. The secondary goal is to extend the results of [30] that were developed for finite-dimensional parameter spaces, to the case where the parameter space is possibly infinite-dimensional.
The first main contribution of this paper is Lemma 3.6. In this result, we identify the generalised eigenpairs of three operator pencils: the first pencil is defined by the Hessian of the negative log-likelihood and the prior precision, the second is defined by the same Hessian and the posterior precision, and the third is defined by the posterior covariance and prior covariance. We use the generalised eigenpairs for the third operator pencil to show that the subspace on which the posterior differs from the prior has dimension equal to the rank of the Hessian of the negative log-likelihood, and that this subspace is spanned by vectors that belong to the Cameron–Martin space of the prior.
The second main contribution of this paper is to formulate and solve some families of Gaussian measure approximation problems, where we measure the error of the approximating Gaussian using either the Kullback–Leibler divergence or the family of Rényi divergences. The first family of measure approximation problems concerns approximating the Gaussian posterior by Gaussians with the true posterior mean but approximate posterior covariances that have the structure of being low-rank negative updates of the prior covariance. We identify optimal solutions for this family in Theorem 4.18. The second family concerns approximating the Gaussian posterior by Gaussians with the true posterior covariance but approximate posterior means. We consider two different types of approximate posterior means and identify optimal solutions for these two types in Theorem 5.6 and Theorem 5.7.
The third main contribution of this paper is to consider the family of measure approximation problems where both the posterior mean and posterior covariance are jointly approximated. In Section 6, we exploit the Pythagorean structure of the formula for the Kullback–Leibler divergence between Gaussians, to show that the optimal solutions for the covariance approximation and mean approximation problems from Theorems 4.18, 5.6 and 5.7 yield optimal solutions for the joint approximation problems.
@@ 1.1 Related literature
The problem of low-rank approximations of covariances of posteriors for linear Gaussian inverse problems posed on finite-dimensional parameter spaces was studied in [14]. In particular, [14, Equation (5)] presents a formula for a low-rank approximation of the posterior covariance that exploits spectral decay in the Hessian of the negative log-likelihood, and [14, Equation (4)] indicates that the error of this low-rank approximation is related to the tail of the spectrum of the so-called ‘prior-preconditioned Hessian’ of the negative log-likelihood.
@@
In [30], a precise formulation of the low-rank posterior covariance approximation problem is given and rigorously analysed, for linear Gaussian inverse problems on finite-dimensional parameter spaces. The low-rank approximation proposed in [14] is shown to be an optimal solution for a family of spectral loss functions that include as special cases the Kullback–Leibler divergence and Hellinger distance between Gaussians with the same mean but different covariances. In addition, the low-rank posterior mean approximation problem was formulated, and two types of posterior mean approximations were studied. The Bayes risk after averaging over realisations of the data was used to quantify the approximation error. For this loss function, optimal solutions to the posterior mean approximation problems were obtained. This approach was further developed for goal-oriented linear Gaussian inverse problems in [29]. Dimension reduction methods for linear Gaussian inverse problems using projections of the data were studied using generalised eigenvalue problems in [15]. The Kullback–Leibler divergence and mutual information were used to quantify the error of the approximating measures.
@@
Dimension reduction for Bayesian inverse problems with possibly nonlinear forward models and non-Gaussian priors appears to have been first analysed in [32]. Joint dimension reduction of parameter and data was studied in [3], also for possibly nonlinear forward models and non-Gaussian priors. The results of [32] were further improved in [21, 20], which derived error bounds in terms of Amari $\alpha$-divergences
@@
$D_{\alpha}(\nu\|\mu)\coloneqq\frac{1}{\alpha(\alpha-1)}\biggr{(}\int\left(\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\right)^{\alpha}\mathrm{d}\mu-1\biggr{)},$
@@
for probability measures $\nu$ and $\mu$ such that $\nu\ll\mu$ and $0<\alpha\leq 1$; see [21, Equation (7)]. The above-cited works consider only the setting of finite-dimensional parameter spaces, and do not consider infinite-dimensional parameter spaces. While [30] provides explicit formulas for the approximation errors, [32, 3, 21, 20] provide only error bounds.
@@
In [25], Kullback–Leibler approximation of probability measures on infinite-dimensional Polish spaces using Gaussians was studied from the calculus of variations perspective. The main results of this work concern existence of minimisers and convergence of a proposed minimisation scheme for identifying the best approximation in a class of approximating Gaussian measures. In our setting, the posterior is already Gaussian, and the approximation classes we consider differ from those in [25].
@@
In [1, Section 3], importance sampling for linear Gaussian inverse problems posed on separable Hilbert spaces is considered. The main result is to identify two types of intrinsic dimension, such that if both dimensions are finite, then absolute continuity of the posterior with respect to the prior holds, and thus importance sampling may be possible. The focus of our work is not to determine whether importance sampling is possible, but rather to identify low-dimensional structure in the Bayesian prior-to-posterior update.
@@ 1.2 Outline
We introduce key notation in Section 1.3 below. In Section 2, we begin by formulating the posterior covariance and posterior mean approximation classes that define our measure approximation problems. In Section 3 we recall the Feldman–Hajek theorem and formulas for the Kullback–Leibler divergence and Rényi divergence of two equivalent Gaussians. We state the first main result of this paper, Lemma 3.6, which identifies the generalised eigenpairs of the three operator pencils mentioned in the introduction and identifies the finite-dimensional subspace on which the posterior differs from the prior. In Section 4, we consider measure approximation problems where the posterior covariance is approximated and identify solutions in Theorem 4.18. In Section 5, we consider the analogous problem for two families of posterior mean approximations, and identify the corresponding solutions in Theorem 5.6 and Theorem 5.7. We then consider joint approximation of the posterior mean and covariance in Section 6. Auxiliary results aare presented in Appendix A, and proofs of the results in this work can be found in Appendix B.
@@ 1.3 Notation
We denote equivalence of two measures $\mu$ and $\nu$ by $\mu\sim\nu$. We write $X\sim\mu$ to denote the fact that the law of a random variable $X$ is $\mu$, and $\mathcal{N}(m,\mathcal{C})$ to denote the Gaussian distribution with mean $m$ and covariance $\mathcal{C}$. The ‘precision’ of $\mathcal{N}(m,\mathcal{C})$ is $\mathcal{C}^{-1}$. We denote the support of a measure $\mu$ by $\operatorname{supp}\mu$. For $p>0$ and $I$ a non-empty interval in $\mathbb{R}$, $\ell^{p}(I)$ denotes the space of $p$-summable sequences, i.e. $\ell^{p}(I)=\{(x_{i})_{i\in\mathbb{N}}\subset I\ :\ \sum_{i\in\mathbb{N}}\lvert x_{i}\rvert^{p}<\infty\}$. We write ‘$a\leftarrow b$’ to denote the replacement of $a$ with $b$. We shall abbreviate ‘orthonormal basis’ by ‘ONB’.
2 Low-rank approximations to linear Gaussian inverse problems

### v3

# Optimal low-rank approximations for linear Gaussian inverse problems on Hilbert spaces Part I: posterior covariance approximation

@@ Abstract
For linear inverse problems with Gaussian priors and Gaussian observation noise, the posterior is Gaussian, with mean and covariance determined by the conditioning formula. Using the Feldman–Hajek theorem, we analyse the prior-to-posterior update and its low-rank approximation for infinite-dimensional Hilbert parameter spaces and finite-dimensional observations. We show that the posterior distribution differs from the prior on a finite-dimensional subspace, and construct low-rank approximations to the posterior covariance, while keeping the mean fixed. Since in infinite dimensions, not all low-rank covariance approximations yield approximate posterior distributions which are equivalent to the posterior and prior distribution, we characterise the low-rank covariance approximations which do yield this equivalence, and their respective inverses, or ‘precisions’. For such approximations, a family of measure approximation problems is solved by identifying the low-rank approximations which are optimal for various losses simultaneously. These loss functions include the family of Rényi divergences, the Amari $\alpha$-divergences for $\alpha\in(0,1)$, the Hellinger metric and the Kullback–Leibler divergence. Our results extend those of Spantini et al. (SIAM J. Sci. Comput. 2015) to Hilbertian parameter spaces, and provide theoretical underpinning for the construction of low-rank approximations of discretised versions of the infinite-dimensional inverse problem, by formulating discretisation independent results.
Keywords: nonparametric linear Bayesian inverse problems, Gaussian measures, low-rank operator approximation, generalised operator eigendecomposition, equivalent measure approximation MSC codes: 28C20, 47A58, 60G15, 62F15, 62G05
@@ 1 Introduction
$m_{\textup{pos}}=m_{\textup{pr}}+\mathcal{C}_{\textup{pos}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}(y-Gm_{\textup{pr}}),\quad\mathcal{C}_{\textup{pos}}=\mathcal{C}_{\textup{pr}}-\mathcal{C}_{\textup{pr}}G^{*}(\mathcal{C}_{\textup{obs}}+G\mathcal{C}_{\textup{pr}}G^{*})^{-1}G\mathcal{C}_{\textup{pr}},$
see e.g. [32, Example 6.23]. It should be noted that $\mathcal{C}_{\textup{pos}}$ does not depend on the realisation $y$ of $Y$.
@@
Computational studies of Bayesian inverse problems on high- but finite-dimensional parameter spaces show that the data is often ‘informative’—i.e., that the posterior differs from the prior—only on a subspace of much lower dimension than the dimension of the parameter space; see e.g. [14]. In [10], a similar subspace is called a ‘likelihood-informed-subspace’. Since the posterior is obtained by reweighting the prior by the likelihood, this likelihood-informed subspace is determined by how the concentration of the likelihood interacts with the concentration of prior. In the case of linear Gaussian inverse problems, the concentration of the likelihood and the concentration of the prior are described by the eigenpairs of the Hessian of the negative log-likelihood and the eigenpairs of the prior precision. These ideas are used in [14] to identify low-rank approximations of the posterior covariance matrix. In [31], the optimality of these posterior covariance approximations with respect to a family of spectral loss functions is shown. In particular, the leading generalised eigenvectors of the Hessian-prior precision matrix pencil build a hierarchy of nested low-dimensional subspaces on which the posterior differs from the prior. If only a few directions in the parameter space need to be stored to be able to approximate the posterior distribution well, then this can be done before observing the data, since these directions are independent of the data. In high but finite-dimensional parameter spaces, this leads to considerable computational and storage savings. Nowadays, the latter is important since read and write operations from memory often form the bottleneck in modern computational hardware, c.f. [25].
So far, the existence of optimal low-rank approximations and likelihood informed subspaces for linear Gaussian prior-to-posterior updates has only been proven for posterior distributions on finite-dimensional parameter spaces. While [32, Example 6.23] provides a formulation of the linear Gaussian inverse problem in infinite dimensions, in the generalisation of the optimal low-rank posterior covariance approximation analysed by [31, Section 2] certain challenges appear.
1.1 Challenges in infinite dimensions
In the finite-dimensional context of [31], the above equation updating $\mathcal{C}_{\textup{pr}}$ to $\mathcal{C}_{\textup{pos}}$ provides a starting point for the approximation procedure. Also the corresponding equation which updates $\mathcal{C}_{\textup{pr}}^{-1}$ to $\mathcal{C}_{\textup{pos}}^{-1}$, c.f. [32, eq. (6.13a)], provides a starting point for the approximation. These are called the ‘prior precision’ and ‘posterior precision’ respectively. An operator pencil involving $\mathcal{C}_{\textup{pr}}^{-1}$ is central in the result of [31, Theorem 2.3]. When the prior distribution is nondegenerate, these can be interpreted as full-rank matrices, that is, finite-dimensional, hence bounded, linear operators. In infinite dimensions, $\mathcal{C}_{\textup{pr}}^{-1}$ and $\mathcal{C}_{\textup{pos}}^{-1}$ are no longer bounded, and they are not even defined on the entire parameter space. In fact, $\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$, the range of the self-adjoint square root of $\mathcal{C}_{\textup{pr}}$, is called the ‘Cameron–Martin space’ of the prior distribution and contains the domain of $\mathcal{C}_{\textup{pr}}^{-1}$. This space is a proper subspace of the parameter space in infinite dimensions, and with probability 1, draws from the prior distribution do not belong to this space. This makes the required analysis of approximations based on the prior-to-posterior precision update more delicate.
Another complication of the infinite-dimensional setting is that, unlike in the finite-dimensional setting, not all approximations of the posterior mean and covariance result in approximate posterior measures that are equivalent to the exact posterior distribution, even if they have the same support. Here, ‘equivalent’ means that the approximate posterior has a density with respect to the exact posterior distribution, and vice versa. Since the prior and the posterior distributions are equivalent for linear Gaussian inverse problems with finite-dimensional data, approximate posteriors which are not equivalent to the exact posterior are also not equivalent to the prior distribution. In fact, nonequivalent Gaussian measures on infinite-dimensional spaces are necessarily mutually singular. That is, they assign full measure to disjoint measurable sets, which is an undesirable property for the approximate posterior and exact posterior/prior distribution to have. Thus, an understanding of which approximate updates of the prior covariance lead to equivalent approximation posterior distributions, with probability 1 with respect to the data $Y$, is needed to construct approximate posterior measures equivalent to the exact posterior.
A third complication is that the analysis of the finite-dimensional setting in [31] relies on certain inherently finite-dimensional results and concepts. For example, in approximating the posterior covariance, a certain loss functional is used to measure the closeness of the approximate posterior covariance to the exact posterior covariance. The coercivity of this loss functional is used to prove some results in the finite-dimensional setting. However, in the infinite-dimensional setting, the analogous coercivity statement does not hold. Also, Fréchet differentiability of this loss functional, useful for finding extreme points of the loss, cannot be deduced in the same way as in the finite-dimensional case, as the latter case relies on the finite-dimensional result of [20, Theorem 1.1].
1.2 Contributions
@@
This work provides a rigorous analysis for the infinite-dimensional version of the Bayesian prior-to-posterior covariance and precision updates and constructs optimal low-rank approximations thereof. We assume a linear Gaussian inverse problem in which the parameter space is a possibly infinite-dimensional separable Hilbert space, the observation space is finite-dimensional and the prior is nondegenerate and has mean zero. We identify optimal Gaussian approximations to the true posterior, keeping the mean fixed, using low-rank measure approximation problems. Our results extend the results of [31] that are developed for finite-dimensional parameter spaces, to the case where the parameter space is possibly infinite-dimensional. This shows a certain dimension independence of the results of [31]. In Part II of this work, [7], we study low-rank posterior mean approximation, and give some insight on joint posterior mean and covariance approximation. We highlight main contributions of this paper.
The first main contribution is Proposition 3.7. In particular:
  • It formulates three operators and their relation in infinite dimensions. The three operators are important in the approximation procedure, and are given by the prior-preconditioned Hessian, the posterior-preconditioned Hessian and the posterior covariance preconditioned with the prior precision. As the relations are one-to-one, these operators contain the same information. It was already noted in previous works in finite dimensions, see e.g. [19, Proposition 10] and [18, Section 3.4.1], that these operators and various other transformations of them contain the same information and are the central object for studying the quality of the finite-dimensional low-rank posterior approximation.
  • It gives one-to-one relations between the above three operators and the Hilbert–Schmidt operator which mixes the prior and posterior covariance in the Feldman–Hajek theorem. The Feldman–Hajek theorem gives necessary and sufficient conditions for equivalence of Gaussian measures, and the connection with the three operators given here, shows that these operators essentially all quantify the amount of similarity and equivalence between the prior and the posterior distribution. This provides intuitive motivation for the importance of this family of operators in the study of optimal posterior approximation.
  • It shows that this family of operators can be diagonalised in the common Cameron–Martin space of the prior and the posterior. In particular, this implies that the diagonalisations of the above family of operators have interpretations as operator pencils as in the finite-dimensional case.
  • It shows that the prior and posterior distribution differ only on a finite-dimensional subspace, which is a subspace of the Cameron–Martin space of both the prior and posterior. Its dimension equals the rank of the Hessian of the negative log-likelihood, or equivalently, the rank of the forward model $G$.
The second main contribution is given by Lemma 4.5 and Proposition 4.6. They are stated for an arbitrary Gaussian measure $\mu_{1}=\mathcal{N}(m_{1},\mathcal{C}_{1})$ with $\mathcal{C}_{1}$ injective. Among the low-rank updates of the covariance $\mathcal{C}_{1}$ and the precision $\mathcal{C}_{1}^{-1}$, these results characterise those low-rank updates which satisfy an equivalence property, namely that when keeping the mean fixed, they correspond to approximate distributions that are equivalent to $\mu_{1}$. Furthermore, these results also characterise the approximate precisions and covariances which correspond to respectively the low-rank covariance and precision updates satisfying this equivalence property. In the Bayesian context, Lemma 4.5 and Proposition 4.6 also show that in infinite dimensions, not all updates of the prior covariance of the form considered in [31] satisfy this equivalence property, and not all updates of the prior covariance that do satisfy this property can be constructed as the inverse of an update of the prior precision considered in [31]. Our results give a necessary and sufficient condition on the range of the low-rank updates, under which such updates of the prior covariance do in fact satisfy the equivalence property.
The third main contribution is to solve a family of Gaussian measure approximation problems in which we approximate the posterior covariance and keep the mean fixed, for example at the exact posterior mean. We consider various loss functions to measure the approximation error of the corresponding approximating Gaussian distribution, including the Rényi divergences, Amari $\alpha$-divergences for $\alpha\in(0,1)$, the Hellinger metric and the forward and reverse Kullback-Leibler divergence. These are all spectral loss functions in the sense that their dependence on the two measures is only via the spectrum of the operators in Proposition 3.7 mentioned above. We ensure that the resulting approximate posterior obtained by approximating the covariance and keeping the mean fixed is equivalent to the exact posterior. Since the posterior covariance and its low-rank approximations are independent of $y$, this equivalence holds for all possible realisations of the data simultaneously. Optimal solutions for the covariance approximation problem and necessary and sufficient conditions for their uniqueness are identified in Theorem 4.21 and Corollary 4.23.
@@ 1.3 Related literature
Low-rank approximation of posterior covariances for linear Gaussian inverse problems posed on finite-dimensional parameter spaces is studied in [14]. In particular, [14, eq. (5)] presents a formula for a low-rank approximation of the posterior covariance that exploits spectral decay in the Hessian of the negative log-likelihood, and [14, eq. (4)] indicates that the error of this low-rank approximation is related to the tail of the spectrum of the prior-preconditioned Hessian of the negative log-likelihood.
@@
In [31], a precise formulation of the low-rank posterior covariance approximation problem is given and rigorously analysed, for linear Gaussian inverse problems on finite-dimensional parameter spaces. The low-rank approximation for the posterior covariance proposed in [14] is shown to be an optimal solution for a family of spectral loss functions that include as special cases the Kullback–Leibler divergence and Hellinger distance between Gaussians with the same mean but different covariances. This approach is further developed for goal-oriented linear Gaussian inverse problems in [30]. Dimension reduction methods for linear Gaussian inverse problems using projections of the data are studied using generalised eigenvalue problems in [15]. The Kullback–Leibler divergence and mutual information are used to quantify the error of the approximating measures.
@@
Dimension reduction for Bayesian inverse problems with possibly nonlinear forward models and non-Gaussian priors appears to have been first analysed in [34]. Joint dimension reduction of parameter and data is studied in [3], for possibly nonlinear forward models and non-Gaussian priors. The results of [34] are further improved in [22, 21], which derived error bounds in terms of Amari $\alpha$-divergences
@@
$D_{\textup{Am}},\alpha(\nu\|\mu)\coloneqq\frac{1}{\alpha(\alpha-1)}\biggr{(}\int\left(\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\right)^{\alpha}\mathrm{d}\mu-1\biggr{)},$
@@
for probability measures $\nu$ and $\mu$ such that $\nu\ll\mu$ and $0<\alpha\leq 1$; see [22, eq. (7)]. The above-cited works consider only the setting of finite-dimensional parameter spaces, and do not consider infinite-dimensional parameter spaces. While [31] provides explicit formulas for the approximation errors, [34, 3, 22, 21] provide only error bounds. In infinite dimensions, [9] proposes a method for sampling the posterior based on the infinite-dimensional likelihood-informed subspace, and identifies the prior-preconditioned Hessian as the fundamental object. However, a rigorous treatment of optimality is not present.
@@
In [26], Kullback–Leibler approximation of probability measures on infinite-dimensional Polish spaces using Gaussians is studied from the calculus of variations perspective. The main results of this work concern existence of minimisers and convergence of a proposed minimisation scheme for identifying the best approximation in a class of approximating Gaussian measures. In our setting, the posterior is already Gaussian, and the approximation classes we consider differ from those in [26].
@@
In [1, Section 3], importance sampling for linear Gaussian inverse problems posed on separable Hilbert spaces is considered. The main result is to identify two types of intrinsic dimension, such that if both dimensions are finite, then absolute continuity of the posterior with respect to the prior holds, and thus importance sampling may be possible. Also [2] considers the setting of linear Gaussian inverse problems on separable Hilbert spaces. In this work, the aim is to analyse the Kullback-Leibler divergence from prior to posterior for optimal experimental design. The focus of our work is not to determine whether importance sampling is possible or study optimal experimental design, but rather to identify low-dimensional structure in the Bayesian prior-to-posterior update.
@@ 1.4 Outline
We introduce key notation in Section 1.5 below. In Section 2, we begin by recalling the infinite-dimensional formulation of the linear Gaussian inverse problem, and formulate the posterior covariance and posterior mean approximation classes that define our measure approximation problems. In Section 3 we recall the Feldman–Hajek theorem, which characterises when two Gaussian measures are equivalent, and recall expressions for the Kullback–Leibler divergence and Rényi divergence of two equivalent Gaussians. We state the first main result of this paper, Proposition 3.7, which identifies the generalised eigenpairs of the three operator pencils mentioned in the introduction and identifies the finite-dimensional subspace on which the posterior differs from the prior. In Section 4, we consider measure approximation problems where the posterior covariance is approximated and identify solutions in Theorem 4.21 and Corollary 4.23. Auxiliary results are presented in Appendix A, and proofs of the results in this work can be found in Appendix B.
@@ 1.5 Notation
For $T\in\mathcal{B}(\mathcal{H},\mathcal{K})$, we denote the adjoint of $T$ by $T^{\ast}\in\mathcal{B}(\mathcal{K},\mathcal{H})$. The space $\mathcal{B}(\mathcal{H})_{\mathbb{R}}$ denotes the space of bounded operators from $\mathcal{H}$ to itself that are additionally self-adjoint. The spaces $\mathcal{B}_{0}(\mathcal{H})_{\mathbb{R}}$, $\mathcal{B}_{00}(\mathcal{H})_{\mathbb{R}}$, $L_{1}(\mathcal{H})_{\mathbb{R}}$ and $L_{2}(\mathcal{H})_{\mathbb{R}}$ and the set $\mathcal{B}_{00,r}(\mathcal{H})_{\mathbb{R}}$ for $r\in\mathbb{N}$ are defined similarly.
For $T\in\mathcal{B}(\mathcal{H})$ we write $T\geq 0$ and $T>0$ if $T$ is nonnegative or positive respectively, i.e. if respectively $\langle Th,h\rangle\geq 0$ or $\langle Th,h\rangle>0$ for all $h\in\mathcal{H}\setminus\{0\}$. If $T\in\mathcal{B}(\mathcal{H})_{\mathbb{R}}$ is nonnegative, then $T^{1/2}$ will denote its nonnegative self-adjoint square root, i.e. $T^{1/2}\in\mathcal{B}(\mathcal{H})_{\mathbb{R}}$. Since $T^{*}T$ is self-adjoint and nonnegative for any $T\in\mathcal{B}(\mathcal{H})$, we may define $\lvert T\rvert\coloneqq(T^{*}T)^{1/2}$.
For $h\in\mathcal{H}$ and $k\in\mathcal{K}$, we interpret the the tensor product $k\otimes h$ as a rank-1 operator in $B(\mathcal{H},\mathcal{K})$, and this operator is $\tilde{h}\mapsto\langle h,\tilde{h}\rangle k$. For $T\in\mathcal{B}_{0}(\mathcal{H,\mathcal{K}})$, $T$ can be written in its ‘singular value decomposition’ (SVD) as a series of rank-1 operators $T=\sum_{i}\sigma_{i}k_{i}\otimes h_{i}$ where $(\sigma_{i})_{i}$ is nonincreasing and nonnegative and $(h_{i})_{i}$ and $(k_{i})_{i}$ are orthonormal sequences in $\mathcal{H}$ and $\mathcal{K}$ respectively, see also Lemma A.5.
A linear operator $T$ from $\mathcal{H}$ to $\mathcal{K}$ which is not necessarily bounded is indicated by $T:\mathcal{H}\rightarrow\mathcal{K}$. Furthermore, $T$ is densely defined if its domain $\operatorname{dom}{T}$ is dense in $\mathcal{H}$. We also write $T:\operatorname{dom}{T}\subset\mathcal{H}\rightarrow\mathcal{K}$ to emphasise the domain of definition of $T$. Thus, $T:\mathcal{H}\rightarrow\mathcal{K}$ generalises the the notion of $T\in\mathcal{B}(H,\mathcal{K})$ in two ways: $\operatorname{dom}{T}$ may be a proper subspace of $\mathcal{H}$ and $T$ need not be bounded on $\operatorname{dom}{T}$. If $T:\mathcal{H}\rightarrow\mathcal{K}$, $S:\mathcal{H}\rightarrow\mathcal{K}$ and $U:\mathcal{K}\rightarrow\mathcal{Z}$ for some separable Hilbert space $\mathcal{Z}$, then $T+S:\mathcal{H}\rightarrow\mathcal{K}$ is defined on $\operatorname{dom}{T}\cap\operatorname{dom}{S}$ and $UT:\mathcal{H}\rightarrow\mathcal{Z}$ is defined on $T^{-1}(\operatorname{dom}{U})$.
Self-adjoint unbounded operators are recalled in Definition A.18 and Definition A.20.
For a densely defined linear operator $S:\operatorname{dom}{S}\subset\mathcal{H}\rightarrow\mathcal{K}$ with domain $\operatorname{dom}{S}\subset\mathcal{H}$, an extension $T$ of $S$ is an operator defined on $\operatorname{dom}{T}\subset\mathcal{H}$, such that $\operatorname{dom}{S}\subset\operatorname{dom}{T}$ and the restriction of $T$ to $\operatorname{dom}{S}$ agrees with $S$. We shall write $S\subset T$ to denote that $T$ is an extension of $S$. If $T$ is bounded, then $T$ is the unique extension of $S$ to all of $\mathcal{H}$.
We let $\Lambda:L_{2}(\mathcal{H})_{\mathbb{R}}\rightarrow\ell^{2}(\mathbb{R})$ be a function that sends a self-adjoint Hilbert–Schmidt operator to its square summable eigenvalue sequence. One possible ordering labels the negative eigenvalues with the even integers and the positive eigenvalues with the odd integers, both ordered decreasingly in absolute value. A different choice is to order the eigenvalues in order of decreasing absolute value. We note that the choice of ordering of $\Lambda$ in two operators $T,S\in L_{2}(\mathcal{H})_{\mathbb{R}}$ is allowed to be different. The precise ordering of eigenvalues that $\Lambda$ assigns to an operator is not important, as we shall only consider compositions of $\Lambda$ with functions on $\ell^{2}(\mathbb{R})$ that are permutation invariant. In analogy to the eigenvalue map $\Lambda:L_{2}(\mathcal{H})_{\mathbb{R}}\rightarrow\ell^{2}(\mathbb{R})$, we define $\Lambda^{m}:L_{2}(\mathcal{Z})_{\mathbb{R}}\rightarrow\mathbb{R}^{m}$ for any $m$-dimensional subspace $\mathcal{Z}\subset\mathcal{H}$ for $m\in\mathbb{N}$ to be the map that sends $X\in L_{2}(\mathcal{Z})_{\mathbb{R}}$ to its eigenvalue sequence, which we order in a nonincreasing way.
We denote equivalence of two measures $\mu$ and $\nu$ by $\mu\sim\nu$. That is, $\mu\sim\nu$ if $\mu$ and $\nu$ are absolutely continuous with respect to each other. The measure $\nu$ is absolutely continuous with respect to $\mu$ if $\mu(A)=0$ implies $\nu(A)=0$ for every measurable set $A$. We denote the support of a measure $\mu$ by $\operatorname{supp}\mu$.
We write $X\sim\mu$ to denote the fact that the distribution of a random variable $X$ is $\mu$. If $X$ has a Gaussian distribution on $\mathcal{H}$, i.e. $\langle X,h\rangle$ is a one-dimensional Gaussian random variable for each $h\in\mathcal{H}$, then we write $X\sim\mathcal{N}(m,\mathcal{C})$, where $m=\mathbb{E}X$ is the mean of $X$ and $\langle\mathcal{C}h,k\rangle=\mathbb{E}\langle h,X-m\rangle\langle X-m,k\rangle$ defines the covariance $\mathcal{C}$ of $X$. The ‘precision’ of $\mathcal{N}(m,\mathcal{C})$ is $\mathcal{C}^{-1}$.
For $I$ a non-empty interval in $\mathbb{R}$, $\ell^{2}(I)$ denotes the space of square-summable sequences, i.e. $\ell^{2}(I)=\{(x_{i})_{i\in\mathbb{N}}\subset I\ :\ \sum_{i\in\mathbb{N}}\lvert x_{i}\rvert^{2}<\infty\}$. If $I\subset\mathbb{R}$ open, then $C^{1}(I)$ denotes the continuously differentiable functions on $I$.
We write ‘$a\leftarrow b$’ to denote the replacement of $a$ with $b$.
2 Low-rank posterior covariance approximations